OpenStudy (anonymous):

LEGEN...waitforit...DARIDDLE: Simple Calculus Question: when is this true: \[\frac{\rm d}{\rm dx} \left(\begin{matrix} \int f(x) \rm dx\end{matrix}\right) \ne \int \left(\begin{matrix} \frac{\rm d}{\rm dx} \rm f(x)\end{matrix}\right)\rm dx \]

hartnn (hartnn):

its always true, unless we define some initial values (for constant of integration) example when f(x) has a constant in it. say, f(x) = x+3 left side = x+3 right side is just x+c if c' = 3, then its becomes false when f(x) does not have a constant in it. say, f(x) = x^2+x left side = x^2+x right side is just x^2+x+c if c=0, then its becomes false

OpenStudy (anonymous):

close enough.

OpenStudy (anonymous):

the answer was actually when f(x) = k where k is a constant...but i suppose it applies to your examples as well